OpenStudy (anonymous):

(WILL GIVE MEDAl) Which of the following points lie in the solution set to the following system of inequalities? y less than or greater to x - 5 y less than or greater to -x - 4 (1, 10) (-1, 10) (10, 1) (1, -10)

terenzreignz (terenzreignz):

Less than or greater to? Or "Less than or equal to"?

terenzreignz (terenzreignz):

\[\Large y \le x-5\]\[\Large y \le-x-4\]

OpenStudy (anonymous):

??? @terenzreignz

terenzreignz (terenzreignz):

Are those^ your inequalities?

OpenStudy (anonymous):

yes @terenzreignz

terenzreignz (terenzreignz):

Well, there's the easier way to do this... Check each point. Say, the point (1,10) Meaning x = 1 and y = 10

terenzreignz (terenzreignz):

And test it out with each inequality to see if it 'fits' It must 'fit' both inequality.

terenzreignz (terenzreignz):

Let's test out the first choice, (1,10) x = 1 y = 10 with the first inequality \[\Large y \le x - 5\]\[\Large \color{blue}{10}\le \color{red}1 - 5\]\[\Large 10 \le -4\] Which clearly isn't true, so the first choice isn't the answer :P Test out the other choices.

OpenStudy (anonymous):

(10, 1)??? @terenzreignz

terenzreignz (terenzreignz):

Show me your work :P

OpenStudy (anonymous):

1 <= 10 - 5 so 1<= 5 @terenzreignz

terenzreignz (terenzreignz):

That's true, but I said it has to fit both inequalities, not just the first. The first point I tested didn't work out with the first one, so I didn't bother checking with the second one. This point, (10, 1) test it with the second inequality...

OpenStudy (anonymous):

1 <= 14 @terenzreignz

terenzreignz (terenzreignz):

Either you made a typo or you didn't notice the negative signs in your second inequality\[\Large y = \color{red}-x \color{red}-4\]

OpenStudy (anonymous):

2 negatives are a positive right?? @terenzreignz

OpenStudy (anonymous):

@terenzreignz